# Multi-Sources Information Fusion Learning for Multi-Points NLOS Localization

### v2

@@ IV-A Datasets
WAIR-D is a mmWave MIMO dataset for researchers to study and evaluate AI algorithms for wireless systems. It uses the ray-tracing method to approximate the Maxwell equations and describes the propagating field as a set of propagating rays, reflecting, diffracting, and scattering over environment elements to obtain the accurate characterization and simulation of electromagnetic propagation. In comparison to the widely-used DeepMIMO dataset, WAIR-D offers a more diverse range of practical propagation environments. This diversity makes it particularly suitable for training robust generalization capabilities in AI models [31].
The parameters employed for generating these datasets are detailed in Tab. I. This comprehensive approach allows us to assess the adaptability of AMDNLoc to different environmental conditions and parameters, demonstrating its potential for wide applications in wireless communications research.
@@
IV-B Visualization of Classification Regions
(a) frequency domain
(b) power, angle and delay domain
(c) combining domain
Figure 4: NLOS pre-classification in multi domains when $\tau_{in}=0.99,\tau_{out}=0.99$
(a) $\tau_{in}=0.98,\tau_{out}=0.99$
(b) $\tau_{in}=0.9,\tau_{out}=0.99$
(c) $\tau_{in}=0.8,\tau_{out}=0.99$
Figure 5: NLOS pre-classification when change $\tau_{in}$ in frequency domain
(a) $\tau_{in}=0.99,\tau_{out}=0.98$
(b) $\tau_{in}=0.99,\tau_{out}=0.95$
(c) $\tau_{in}=0.99,\tau_{out}=0.9$
Figure 6: NLOS pre-classification when change $\tau_{out}$ in frequency domain
The division of NLOS regions within AMDNLoc is primarily guided by the termination thresholds, $\tau_{in}$ and $\tau_{out}$, in the PFCFR matched filter. Initially, we visualize the pre-classification regions across multiple domains with $\tau_{in}=\tau_{out}=0.99$, as depicted in Fig. 4. Notably, Fig. 3(c) demonstrates more flexible and reasonable boundaries, attributable to the deep fusion of features in combining domains.
Subsequently, to assess the impact of varying termination thresholds on the localization, we alter $\tau_{in}$ and $\tau_{out}$ independently, as illustrated in Fig. 5 and Fig. 6. We observe that changes in $\tau_{out}$ significantly affect the number of categories more than that in $\tau_{in}$. This is because $\tau_{out}$ constrains the inter-category similarity, and a lower value tends to merge more categories. Hence, it acts as a coarse tuning threshold and should be set to a higher value. Conversely, $\tau_{in}$ governs the intra-category similarity, influencing the classification of fingerprints into categories. Adjusting $\tau_{in}$ therefore has a fine tuning effect on the region division.
To determine optimal values for $\tau_{out}$ and $\tau_{in}$, we compute the MSE by varying $\tau_{in}$ while keeping $\tau_{out}$ constant at 0.99, as shown in Tab. II. Additionally, we vary the number of data cleansing samples to 2 or 10. We find that decreasing $\tau_{in}$ reduces MSE to a certain extent. Furthermore, removing more special categories leads to smaller scene coverage but higher localization accuracy. Although a trade-off exists between coverage area and MSE, opting for a lower coverage rate can enhance localization accuracy. This approach aligns with our expectations, considering that if a MT encounters a special sample, it can easily move a short distance to achieve more accurate position estimation.
In conclusion, our pre-classification method is highly effective and reliable. It divides a large outdoor area of $250\times 250$ meters, with numerous buildings, into small regions with irregular boundaries. Each region exhibits similar fingerprint distribution characteristics, and there is a clear distinction between regions, significantly mitigating the issue of fingerprint distribution heterogeneity. These characteristics make our approach applicable in various contexts such as channel estimation [32, 33], channel extrapolation [34], CSI design for sensing [35, 36], and beamforming [37, 38].
Table II: The MSE of changing the $\tau_{in}$ when $\tau_{out}=0.99$.
$\tau_{in}$ Data Cleansing Sample Number Region Covering Rate MSE(m)
0.99 2 0.984 2.54
10 0.918 2.23
0.98 2 0.964 2.66
10 0.823 2.28
0.9 2 0.967 2.62
10 0.782 2.57
0.8 2 0.976 2.36
10 0.824 2.17
@@ IV-C Effectiveness of Classification Methods
IV-C1 Comparison of Multiple Processing Results (MPR)
In this part, we compare our proposed algorithm with several baseline algorithms for NLOS environments using a ULA. To facilitate a fair comparison, we uniformly employ ResNet18 as the backbone for feature extraction in all methods. The feature extractors for CFR and ADCAM in different comparison methods share the same structure to maintain consistency. The structures of all AI-based baselines have been carefully optimized through cross validation to ensure fairness in comparison. The baseline algorithms include:
@@ IV-C1 Comparison of Multiple Processing Results (MPR)
    Res_CFR/Res_ADCAM/Res_CFRADCAM: These methods utilize either CFR, ADCAM, or a combination of both as input for the feature extractor before feeding the features into a linear classifier to predict coordinates.
@@
    Res_multi_CFRADCAM/Res_multi_CFRperfectADCAM: Building on Res_CFRADCAM, these baselines input the obtained features into multiple linear classifiers for the coordinate prediction. Specifically, Res_multi_CFRperfectADCAM replaces ADCAM with additional inputs such as AOA, AOD, distance, and gain from the first arrival path.
Figure 7: MSE with epochs of MPR.
Figure 8: CDF curves of MPR.
Fig. 7 provides convincing evidence of our methodology’s effectiveness. Initially, we observe that the MSE for Res_ADCAM decreases rapidly with an increase in epochs but eventually stabilizes near the higher MSE of Res_CFR. This pattern emerges because, while ADCAM is sparser than CFR, it does not contain as much information, so using complex network structures is more prone to overfitting. Then, we notice a significant improvement when combining CFR and ADCAM as input data. By preprocessing and fusing these inputs, and employing a segment-specific linear classification mechanism, our results show a remarkable improvement, reducing the error from 9 meters to 2.54 meters in just 150 epochs. This improvement is attributed to AMDNLoc’s focus on the variations between locations and fingerprints with similar data distributions, allowing for a more accurate reflection of the true nature of the localization-fingerprint relationship.
Moreover, Fig. 8 underscores this advancement, revealing that over 60% of positioning errors in our model fall within a 2-meter range. In contrast, the baseline models only demonstrate a 20% probability of achieving such accuracy. These findings unequivocally affirm the superiority of our integrated approach in enhancing the accuracy and reliability of positioning in MIMO-OFDM systems.
IV-C2 Comparison of Pre-processing on CFR or ADCAM
In this part, we evaluate the performance of CFR and ADCAM as inputs for feature extraction and localization, both individually and in conjunction with our pre-classification technique as illustrated in Fig. 9. The approaches are as follows:
  1. 1.
    CFR/ADCAM_direct: This method involves using either CFR or ADCAM as the input for the feature extractor. The extracted features are then processed through a linear classifier to predict coordinates.
  2. 2.
    CFR/ADCAM_ours: This approach employ multiple linear classifiers for coordinate prediction building upon CFR/ADCAM_direct. The number of classifiers corresponds to the number of pre-classifications in the frequency or power, angle, and delay domains.
The results, as depicted in Fig. 9, reveal some key insights. Firstly, we observe that using ADCAM alone as a fingerprint for coordinate prediction yields better results than using CFR alone, which aligns with the existing results in the [12]. Secondly, upon applying our PFCFR matched filter, the performance of CFR significantly improves, eventually surpassing that of ADCAM as the iterations progress. Similarly, a substantial enhancement is noted in the use of ADCAM after iterative clustering. These findings indicate that our pre-classification method, combined with an array of linear classifiers, is highly effective and valuable. This holds true not only for the combined use of CFR and ADCAM, but also when each is used independently.
Figure 9: MSE with epochs on CFR and ADCAM.
IV-D Impact of System Parameters
In this subsection, the impact of various system parameters are investigated to validate the superiority and universality of the proposed approach.
IV-D1 Antenna Array
Figure 10: MSE with epochs of MPR with UPA array.
Figure 11: CDF curves of MPR with UPA array.
We alter the transmitting antenna from ULA with a configuration of [1,64,1] to UPA of [1,8,8]. The outcomes of this change are depicted in Fig. 10 and Fig. 11. Upon comparison with the results from the ULA array shown in Fig. 8 and Fig. 9, we observe that the trend and relative positioning of each curve do not exhibit significant differences. However, it is noteworthy that, despite the total number of transmitting antennas remains the same, the planar array configuration yields better results in terms of MPR. With the implementation of AMDNLoc, the positioning error decreases from 6.95 meters with CFR to 1.46 meters, which constitutes a remarkable 79% enhancement in the localization accuracy for outdoor multi-point NLOS scenarios.
IV-D2 Signal to Noise Ratio (SNR)
To analyze the impact of SNR on the localization accuracy, we first normalize the energy of the transmitted signal and calculate the average energy of the channel. Then, based on the SNR, we determine the variance of the noise and generated a complex Gaussian matrix. This matrix is incorporated into the fingerprint to observe the variation in the optimal MSE with varying SNR levels, as illustrated in Fig. 12.
Our findings show that although the accuracy of localization of MPR generally improves with an increase in SNR, the variation in the results of AMDNLoc with respect to SNR is comparatively minimal. Notably, in 10-15dB, AMDNLoc demonstrates localization estimation results that are comparable to those achieved at higher SNR levels and significantly better than those obtained without using AMDNLoc. This performance indicates that AMDNLoc is more robust and better suited to the diverse conditions of real-world scenarios, especially where SNR may vary.
IV-D3 Maxpathnum
Maxpathnum, representing the maximum number of paths between a transmitter and a receiver, is a crucial parameter in ray-tracing software. A higher Maxpathnum value implies more scattering and refraction, leading to increased complexity in the multi-path effect. As shown in Fig. 13, when Maxpathnum is incrementally increased from 10 to 30, there is a noticeable degradation in the optimal MSE of MPR, reflecting the intensified multi-path effects.
In contrast, the localization error of AMDNLoc in relation to Maxpathnum remains smaller compared to other methods. This can be attributed to the deep fusion of features across multiple domains within AMDNLoc. Such fusion ensures that variations in the fingerprint dataset do not significantly impact the overall classification regions. Consequently, our framework exhibits greater stability and adaptability, particularly in static environments with dynamic changes. This ability to maintain accuracy despite increasing Maxpathnum underscores the robustness of AMDNLoc in complex multi-path scenarios.
Figure 12: MSE with SNR of MPR.
Figure 13: MSE with Maxpathnum of MPR.
IV-E Generalization Ability
IV-E1 Multiple Scenarios and Datasets
We conduct a comprehensive validation of the AMDNLoc framework across various scenarios generated through ray tracing, utilizing different datasets. The selected scenarios include DeepMIMO O1 blockage, WAIR-D 00247, and WAIR-D 00743. The results, compiled after 400 epochs, are presented in Tab. III. A key observation is that the segment-specific linear array demonstrates greater improvements in accuracy, especially in scenarios with an increased number of buildings compared to those without the array.
Furthermore, in datasets with a higher density of MTs, such as the DeepMIMO O1 blockage scenario where the spatial distribution of sample MTs is more uniform and compact, the use of ADCAM as fingerprints yields results remarkably close to those obtained using AOA, AOD, distance, and gain. Notably, despite the augmented complexity of these datasets, the optimal MSE consistently remains within a 2-meter range. This consistent performance, even under varying conditions and increased dataset complexity, highlights the robustness and adaptability of the proposed AMDNLoc framework across diverse scenarios and MT spatial distributions.
Table III: MSE of multiple scenarios and datasets with UPA array.
Methods DeepMIMO O1 blockage WAIR-D 00247 WAIR-D 00743
Res_CFRADP 1.66 2.23 4.21
Res_multi_CFRADP 1.11 1.54 1.99
Res_multi_CFRperfectADP 0.92 1.11 1.25
IV-E2 Multiple Feature Extractors
Table IV: MSE and running time of different feature extractors.
Epochs Items DenseNet121 ResNet18 EfficientformerV2
150 MSE(m) 2.44 2.54 2.58
Time(s) 5988 2192 8050
250 MSE(m) 2.27 2.24 2.28
Time(s) 9946 3655 13415
We conduct an evaluation to assess the adaptability of the AMDNLoc framework with various neural networks, using different feature extractors such as DenseNet121 and EfficientformerV2, as outlined in Tab. IV. Through this comparative analysis, we found that ResNet18 offers the most effective balance between time efficiency and accuracy. DenseNet18, while achieving better accuracy, is slower than ResNet18 due to its use of dense connections instead of skip connections.
While lightweight Vision Transformer (ViT) models are generally known for outperforming ResNet models in certain applications [39], in our specific context, ResNet18 demonstrates superior speed and performance compared to the latest lightweight ViT models. This can be attributed to the fact that DCNN has historically been more adept at extracting features from CSI data [40]. Our pre-classification function effectively organizes the data, ensuring similar fingerprint distributions within each class, which is conducive to feature extraction by ResNet. However, the primary challenge with ViT lies in handling inductive bias. In theory, our pre-classification should provide a favorable inductive bias for ViT, potentially leading to better results. Nonetheless, ViT’s approach of decomposing images into batches for network training poses a limitation. Given the high similarity within each batch of CFR data, ViT struggles to effectively manage this translation and similarity, leading to suboptimal outcomes. Despite this, we believe that further exploration and modifications of the existing ViT model could yield improvements, presenting a promising avenue for future research.
IV-E3 From Only NLOS to NLOS and LOS
We have expanded our study to include scenarios that encompass both NLOS and LOS conditions. The classification areas under these combined conditions are visualized in Fig. 14. Similar to the observations in Fig. 3(c), our AMDNLoc framework demonstrates effective intra-group continuity and inter-group discrimination. Notably, it achieves an accuracy of less than 2 meters.
It is important to highlight that CSI is different in NLOS and LOS scenarios. Consequently, existing research mainly focuses on either mitigating or identifying NLOS conditions prior to localization [41, 42]. However, we present a novel perspective by integrating both NLOS and LOS conditions to a certain extent. The effectiveness of AMDNLoc in capturing the distinct feature distributions of channel information suggests its broader applicability that is not limited to multi-point fingerprint localization scenarios.
Figure 14: NLOS and LOS pre-classification in combining domain.
@@ V Conclusion
In the paper, we proposed a novel multi-sources information fusion learning framework that makes full use of the multi-path features across the frequency, power, angle and delay domain as fingerprints to tackle the inherent heterogeneity issue of fingerprint distributions. We explored a two-stage matched filter for PFCFR corresponding to the distribution of CFR, and fused it with ADCAM classification region after centroid-based clustering method. What’s more, a segment-specific linear classifier mechanism sharing the same feature extractor was utilized to build regression relationship between fingerprint and locations after eliminating the negative effect of the region covariant. Our numerical experiments demonstrated that AMDNLoc achieves SOTA results, outperforming traditional convolutional neural networks in datasets such as WAIR-D and DeepMIMO. These findings highlighted AMDNLoc’s exceptional accuracy, interpretability, adaptability, and scalability, marking a significant advancement in the field of wireless communication and localization.

### v3

@@ IV-A Datasets
WAIR-D is a mmWave MIMO dataset for researchers to study and evaluate AI algorithms for wireless systems. It utilizes ray-tracing method, which is based on a high-frequency approximation to the Maxwell equations and describes the propagating field as a set of propagating rays, reflecting, diffracting, and scattering over environment elements to obtain the accurate characterization and simulation of electromagnetic propagation. The parameters used to generate the dataset are listed in Tab. I.
@@
IV-B Visualization of Classification Areas
Firstly, we visualize the classification results of the scene as shown in Fig. 4. It can be seen that our pre classification has excellent effectiveness, dividing an outdoor scene with a large area of $250\times 250$ meters and many buildings into lots of small areas with irregular boundaries. The data within each region has relatively similar distribution characteristics, and there is good distinguishability between regions as well, greatly reducing the data distribution heterogeneity problem. By leveraging these features, it can be used for channel estimation, channel inference, beamforming and so on.
@@ IV-C Effectiveness of Classification Methods
We compare the proposed algorithm with the following baseline algorithms for NLOS when the antenna array is ULA.
@@
    Res_CFR/Res_ADCAM/Res_CFRADCAM: Use CFR or ADCAM or CFR and ADCAM together as input for the feature extractor, and then put the obtained features into a linear classifier to obtain the predicted coordinates.
@@
    Res_multi_CFRADCAM/Res_multi_CFRperfectADCAM: Compared to Res_CFRADCAM, these two baselines put the obtained features into multiple linear classifiers to obtain the predicted coordinates. Res_multi_CFRperfectADCAM replaces ADCAM with AOA, AOD, distance, gain of the first arrival path as input.
Figure 5: MSE with epochs of multiple processing results.
Figure 6: CDF curves of multiple processing results.
Fig. 5 presents compelling evidence of the efficacy of our methodology. Firstly, as epochs increase, the MSE for Res_ADCAM rapidly decreases, but ultimately close to that of Res_CFR which is the highest. However, by adopting our AMDNLoc which fuses ADCAM and CFR deeply, the optimal MSE can be improved from 6.34 to 2.14 meters. This is because ADCAM is inherently sparser than CFR, and relying solely on either fingerprint can’t fully represent the multi-path features. What’s more, AMDNLoc is more focused on the changes between the locations and the fingerprints with the similar data distribution, so it can better reflect the true nature of the location-fingerprint relationship. Further supporting our findings, Fig. 6 illustrates that over 60% of positioning errors in our model fall within a 2-meter range. This is a significant increase compared to the baseline models, which only demonstrate a 20% probability of positioning errors within the same range. These results unequivocally demonstrate the superiority of our integrated approach in achieving more accurate and reliable positioning in MIMO-OFDM systems.
@@ V Conclusion
In the paper, we proposed a novel multi-sources information fusion learning framework named AMDNLoc that makes full use of the multi-path feature across the frequency, power, angle and delay domain as fingerprints to tackle the inherent heterogeneity issue of fingerprint distributions. We explore a two-stage matched filter for PFCFR corresponding to the distribution of CFR, and fuse it with ADCAM classification region after centroid-based clustering method. What’s more, a segment-specific linear classifier mechanism sharing the same feature extractor is utilized to build regression relationship between fingerprint and locations after eliminating the negative effect of the region covariant. Numerical experiments have shown that the AMDNLoc achieves SOTA results compared with traditional convolutional neural networks on the WAIR-D.